# Reinforcement learning-based statistical search strategy for an axion model from flavor

### v1

@@ 1 Introduction
The machine learning-based search strategy is of particular attention not only for collider physics but also for particle phenomenology. So far, the quantitative evaluation of physical observables relied on various optimization methods to search for desirable parameters in a model and to predict new phenomena at high-energy scales. However, the analysis often requires computational costs to exhaust huge parameter spaces in the model. Therefore, the machine learning-based analysis will be expected as a powerful approach to searching for physics beyond the Standard Model and reducing the computational costs. Among the machine learning techniques, we focus on the reinforcement learning (RL). In contrast to supervised learning and unsupervised learning, the RL allows us to find solutions autonomously from a small amount of data.
@@
In this paper, we propose a new approach to search for physics beyond the Standard Model, i.e., a RL-based search strategy. In particular, we apply the RL-based search strategy to the minimal axion model with a global $U(1)$ flavor symmetry proposed in Refs. Ema:2016ops ; Calibbi:2016hwq , as a fascinating example solving the flavor and cosmological puzzles in the Standard Model, i.e., mass hierarchies and mixing angles of quarks and leptons, an origin of dark matter, and the inflation. In this model, the flavor structure of quarks and leptons can be explained by using the Froggatt-Nielsen (FN) mechanism Froggatt:1978nt . Since the $U(1)$ flavor symmetry is identified with the Peccei-Quinn (PQ) symmetry, the QCD axion originated from the Nambu-Goldstone (NG) boson of the spontaneously broken $U(1)$, so-called a flaxion field, can solve the strong CP problem via the PQ mechanism PhysRevLett.38.1440 ; PhysRevD.16.1791 . Then, a proper value of the flaxion decay constant also explains the dark matter abundance in our universe. Furthermore, the successful inflation mechanism can be realized by the radial component of the $U(1)$ breaking field with the proper Lagrangian.
@@
So far, the flavor structure of quarks and leptons was studied by using the RL-based method in Ref. Nishimura:2023nre . In particular, agents of RL succeeded in finding the best solutions to be consistent with the mass hierarchies and the mixing angles of quarks and leptons. This work motivates us to apply the RL-based strategy to explore the cosmological aspects of the minimal axion model with the global $U(1)$ flavor symmetry. Since couplings between the flaxion and quarks/leptons are determined by their $U(1)$ charges, the flaxion-photon coupling as well as the flaxion decay constant is sensitive to $U(1)$ charge assignment. We utilize the RL-based strategy to find the best $U(1)$ charges of quarks/leptons solving the flavor and cosmological puzzles in the Standard Model. By comparing the traditional optimization methods with the RL-based search strategy, we argue that the RL-based search method for discrete parameters is faster than conventional computational methods. Since the predicted flavor structure will be completely different for the choice of $U(1)$ charge assignment, the evaluation function of the model rapidly fluctuates under the change of discrete parameters. In such a situation, the traditional numerical optimization with the gradient descent method is difficult to apply to this system, and the RL-strategy for discrete parameters has an advantage over the conventional methods.
@@
This paper is organized as follows. After briefly reviewing the flaxion model and constraints from flavor physics and cosmology in Sec. 2, we organize the RL-based search strategy in Sec. 3. In Sec. 4, we verify how fast the RL-based method finds the proper $U(1)$ charge assignments solving the problems in the Standard Model. In Sec. 5, we show training results and discuss the phenomenology of the flaxion with an emphasis on the flaxion-photon coupling which will be explored in the future experiments such as DMRadio-$\mathrm{m}^{3}$ DMRadio:2022pkf . Sec. 6 is devoted to conclusion and discussion. In Appendix A, we list our finding $U(1)$ charge assignments of the leptons.
@@ 2.1 Model
$-{\cal L}\supset~{}$ $y^{u}_{ij}\left(\frac{\phi}{M}\right)^{n^{u}_{ij}}\bar{Q}_{i}H^{c}u_{j}+y^{d}_{ij}\left(\frac{\phi}{M}\right)^{n^{d}_{ij}}\bar{Q}_{i}Hd_{j}+y^{l}_{ij}\left(\frac{\phi}{M}\right)^{n^{l}_{ij}}\bar{L}_{i}Hl_{j}$
@@
where $\{Q_{i},u_{i},d_{i},L_{i},l_{i},N_{i},H\}$ denote the left-handed quarks, the right-handed up-type quarks, the right-handed down-type quarks, the left-handed leptons, the right-handed charged leptons, the right-handed neutrinos, and the SM Higgs doublet with $H^{c}=i\sigma_{2}H^{\ast}$, respectively. We assume that a mass of the right-handed neutrinos $M$ is corresponding to a cut-off scale of the model to explain the light neutrino masses through the seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia . The Yukawa couplings $\{y^{u}_{ij},y^{d}_{ij},y^{l}_{ij},y^{\nu}_{ij},y^{N}_{ij}\}$ are ${\cal O}(1)$ real constants.
@@
$\begin{split}V_{\rm CKM}&=U^{u\dagger}U^{d}\\ &=\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-i\delta_{\rm CP}}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}e^{i\delta_{\text{CP}}}&c_{12}c_{23}-s_{12}s_{23}s_{13}e^{i\delta_{\text{CP}}}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}e^{i\delta_{\text{CP}}}&-c_{12}s_{23}-s_{12}c_{23}s_{13}e^{i\delta_{\text{CP}}}&c_{23}c_{13}\end{pmatrix},\end{split}$ (2.10)
@@
$f_{a}\equiv\frac{\sqrt{2}v_{\phi}}{N_{\mathrm{DW}}}=\frac{\sqrt{2}\eta M}{N_{\mathrm{DW}}}\,,$ (2.20)
@@
For phenomenological study, it is also useful to consider an effective coupling between the flaxion and the photon. The effective Lagrangian of the interaction is given by
@@ 2.2 Constraints from flavor physics
Comparing with a current experimental bound from NA62 collaboration NA62:2020pwi ; NA62:2021zjw , Br$(K^{+}\to\pi^{+}a)\lesssim(3-6)\times 10^{-11}$ at $90\%$ C.L., a lower bound on $f_{a}$ is given by
$f_{a}\gtrsim\frac{7.6\times 10^{11}\,\mbox{GeV}}{N_{\mathrm{DW}}}\left|\frac{(\kappa^{d}_{\rm AH})_{12}}{m_{s}-m_{d}}\right|.$ (2.27)
This bound can be evaluated by determining appropriate parameters of the FN model.
@@ 2.3 Constraints from cosmology
As discussed later, the misalignment angle is fixed through Eq. (2.29) for each model found by RL, so the upper bounds of $H_{\rm inf}$ is given for each of the models.
@@ 3 Reinforcement learning
In RL, the subject of learning is called an agent, and the problem to be solved is called an environment. The learning process is constructed from three stages: observing the environment, choosing an action, and getting rewards. By repeating these procedures, it is known that the agent autonomously acquires a principle of actions that maximizes the sum of rewards. Details of RL and Q-learning, which are basic ingredients in RL, are explained in Ref. RL . In this paper, we adopt Deep Q-Network (DQN) as a RL algorithm, and use the package “gym” developed by OpenAI. Since the architecture in this work is similar to our previous work Nishimura:2023nre , we briefly review the RL, and setups of the learning will be explained by focusing on the key points. For more details about the architecture, see Ref. Nishimura:2023nre .
The flavor structure of quarks and leptons is determined by the charge shown in Eq. (2.2). In Ref. Nishimura:2023nre , many realistic models of the lepton sector are found in contrast to the quark sector. In other words, there are many varieties of the $U(1)$ charges of leptons with successful mass hierarchies and flavor structure in the quark sector. Hence, at the moment, we will perform RL with an emphasis on the quark sector.
@@ 3.1 The environment
with $i=1,2,3$, so the vector has 10 elements. This is an input data for neural networks which decide an action of the agent. To avoid a hierarchy of $U(1)$ charges, we take the value of each charge within:
@@
3.2 Neural Network
We utilize neural networks to determine an action of the agent. Their architecture is summarized in Table 1. The activation function is chosen as the SELU function for hidden layers and the softmax function for an output layer. We employ the ADAM optimizer in TensorFlow DBLP:journals/corr/AbadiABBCCCDDDG16 , and the Huber function is used as a loss function. Hyper-parameters are set as $N_{\rm ep}=10^{5}$ for an episode number and $N_{\rm step}=32$ for a step number. Moreover, a batch size is 32, an epoch number is 32, and a learning rate is $\alpha=2.5\times 10^{-4}$, respectively  [^1] For simplicity, we use the same hyper-parameters as in the previous study, but the choice of hyper-parameters is also important as a factor that determines the success or failure of the learning. There are various optimization methods for hyper-parameters. Then, “Hyperopt” Bergstra_2015 and “Optuna” 10.1145/3292500.3330701 are known as libraries that provide tuning methods based on Bayesian optimization. In general, a “black box function” refers to a function whose functional form cannot be revealed in advance and whose value is determined only after the arguments are given. Bayesian optimization and genetic algorithms are effective methods for solving optimization problems for black box functions..
@@ 3.3 Agent
where $A$ corresponds to $\{Q_{i},u_{i},d_{i},\phi\}$. At an initial stage of the learning, for each environment, the ${\cal O}(1)$ coefficients in Yukawa terms Eq. (2.1) are picked up from the two Gaussian distribution with an average ±$\pm 1$ and a standard deviation 0.25. After the training with the neural networks introduced in the previous subsection, the coefficients are optimized to proper values by the Monte-Carlo simulation. Note that the coefficient $y^{u}_{33}$ was fixed at 0.4 in both the initialization and the optimization to ensure that the mass of the top quark is reproduced. In addition, the agent chooses an action based on $\varepsilon$-greedy method. Following the same way as Ref. Nishimura:2023nre , we adopt $\varepsilon_{0}=1,\ r=0.99999$ and $\varepsilon_{\rm min}=0.01$.
@@
    For the mixing angles, we focused on the reproduction of the feature of weak mixing rather than exact values that take the effect of renormalization into account. In addition, only one flavon field is introduced in the Lagrangian Eq. (2.1). Thus, $\mathcal{C}$ does not include the evaluation of the CP phase.
@@
A large intrinsic value indicates that an obtained charge assignment well reproduces the renormalized masses and mixings  [^2] In the present study, the intrinsic value is defined the same as in Ref. Nishimura:2023nre to be based on the findings of the previous study. If the $\chi^{2}$ value is adopted for intrinsic value, the search can be conducted with more statistical significance.. Such charge assignment is called a terminal state. Specifically, the terminal state is defined to satisfy following requirements:
@@
In this paper, we adopt $V_{0}=10.0$, $V_{1}=1.75$, and $V_{2}=0.2$. Here, $V_{1}=1.75$ ($V_{2}=0.2)$ means that a ratio of the calculated masses (mixings) to the renormalized masses (observed mixings) satisfies $1.78\times 10^{-2}\leq r_{\rm mass}\leq 56.2$ ($0.63\leq r_{\rm mixings}\leq 1.58$).
Note that we change the threshold value from $V_{1}=1.0$ used in Ref. Nishimura:2023nre to $V_{1}=1.75$. In that work, experimental masses at the electroweak scale, which are shown in Particle Data Group ParticleDataGroup:2022pth , were used. On the other hand, this study considers renormalized masses that are smaller than those used in the previous research. For example, the up-quark mass shown in Ref. ParticleDataGroup:2024prd is $m_{u,\mathrm{exp}}=2.16\,\mbox{MeV}$, and an allowed range for a terminal state in the case of $V_{1}=1.0$ is $0.216\,\mbox{MeV}\leq m_{u}\leq 21.6\,\mbox{MeV}$. However, $m_{u,\mathrm{RG}}=0.49\,\mbox{MeV}$ for $M=10^{14}\,\mbox{GeV}$ as shown in Table 2, so an allowed range with the same $V_{1}$ is $0.049\,\mbox{MeV}\leq m_{u}\leq 4.9\,\mbox{MeV}$. In other words, a tolerance $\Delta m_{u}$ to reach a terminal state varies in the following way:
@@
Since changes of the renormalized masses from $10^{14}\,\mbox{GeV}$ to $10^{17}\,\mbox{GeV}$ are relatively small, the difference of tolerance has little effect on search results at each scale. Then, after optimizing the ${\cal O}(1)$ coefficients, we manually extract terminal states where $n_{ij}$ is positive. Moreover, at this time, the reproducibility of the masses is ensured by extracting models that satisfy $V_{1}=1.75\,(0.1\leq r_{\rm mass}\leq 10.0)$.
Let us denote the charge assignment ${\cal Q}$ as observed by the agent, and ${\cal Q}^{\prime}$ followed by the action $\mathfrak{a}$. For a pair of the assignment and the action $({\cal Q},\mathfrak{a})$, we give the agent rewards $\mathcal{R}$. The rewarding procedure is exactly the same as in Ref. Nishimura:2023nre ; thereby we omit a detailed description in this paper  [^3] In many cases, the appropriate design of rewards requires heuristic construction depending on the environment and action space. In contrast, it has been pointed out that the design may be systematized by combining Bayesian statistics with RL levine2018reinforcementlearningcontrolprobabilistic . Alternatively, by learning based on models in Bayesian statistics, various algorithms have been developed that can perform learning even with small amounts of data. “PILCO” 10.5555/3104482.3104541 is one example.. We adopt $\mathcal{R}_{\rm offset}=-10$ and $\mathcal{R}_{\rm term}=100$.

### v2

@@ 1 Introduction
The machine learning-based search strategy is of particular attention not only for collider physics but also for particle phenomenology. So far, the quantitative evaluation of physical observables has relied on various optimization methods to search for desirable parameters in a model and to predict new phenomena at high-energy scales. However, this analysis often requires computational costs due to the need to explore vast parameter spaces in the model. Therefore, the machine learning-based analysis is anticipated to a powerful approach for investigating physics beyond the Standard Model and reducing the computational costs. Among the machine learning techniques, we focus on the reinforcement learning (RL). In contrast to supervised learning and unsupervised learning, the RL enables us to autonomously discover solutions using a limited amount of data.
@@
In this paper, we propose a new approach to search for physics beyond the Standard Model, i.e., a RL-based search strategy. In particular, we apply the RL-based search strategy to the minimal axion model with a global $U(1)$ flavor symmetry proposed in Refs. Ema:2016ops ; Calibbi:2016hwq  [^1] See, Refs. Davidson:1981zd ; Wilczek:1982rv for earlier works., as a fascinating example of solving the flavor and cosmological puzzles in the Standard Model, i.e., mass hierarchies and mixing angles of quarks and leptons, an origin of dark matter, and the inflation. In this model, the flavor structure of quarks and leptons can be explained by using the Froggatt-Nielsen (FN) mechanism Froggatt:1978nt . Since the $U(1)$ flavor symmetry is identified with the Peccei-Quinn (PQ) symmetry, the QCD axion originated from the Nambu-Goldstone (NG) boson of the spontaneously broken $U(1)$, so-called a flaxion field, can solve the strong CP problem via the PQ mechanism PhysRevLett.38.1440 ; PhysRevD.16.1791 . Then, a proper value of the flaxion decay constant also explains the dark matter abundance in our universe. Furthermore, the successful inflation mechanism can be realized by the radial component of the $U(1)$ breaking field with the proper Lagrangian.
@@
So far, the flavor structure of quarks and leptons was studied using the RL-based method in Ref. Nishimura:2023nre . In particular, agents of RL succeeded in finding the best solutions to be consistent with the mass hierarchies and the mixing angles of quarks and leptons. This work motivates us to apply the RL-based strategy to explore the cosmological aspects of the minimal axion model with the global $U(1)$ flavor symmetry. Since couplings between the flaxion and quarks/leptons are determined by their $U(1)$ charges, the flaxion-photon coupling as well as the flaxion decay constant are sensitive to the assignment of these charges. We utilize the RL-based strategy to find the best $U(1)$ charges of quarks/leptons solving the flavor and cosmological puzzles in the Standard Model. By comparing traditional optimization methods with the RL-based search strategy, we argue that the RL-based approach for discrete parameters is more efficient than the conventional techniques. Since the predicted flavor structure varies significantly with different $U(1)$ charge assignments, the evaluation function of the model exhibits rapid fluctuations in response to changes in discrete parameters. In such a situation, the traditional numerical optimization with the gradient descent method is difficult to apply to this system, and the RL-strategy for discrete parameters has an advantage over the conventional methods.
@@
This paper is organized as follows. After briefly reviewing the flaxion model and constraints from flavor physics and cosmology in Sec. 2, we organize the RL-based search strategy in Sec. 3. In Sec. 4, we verify how fast the RL-based method finds the proper $U(1)$ charge assignments solving the problems in the Standard Model. In Sec. 5, we show training results and discuss the phenomenology of the flaxion with an emphasis on the flaxion-photon coupling which will be explored in the future experiments such as $\mathrm{m}^{3} DMRadio$:2022pkf . Sec. 6 is devoted to conclusion and discussion. In Appendix A, we list our finding $U(1)$ charge assignments of the leptons.
@@ 2.1 Model
$+y^{\nu}_{ij}\left(\frac{\phi}{M}\right)^{n^{\nu}_{ij}}\bar{L}_{i}H^{c}N_{j}+\frac{y^{N}_{ij}}{2}\left(\frac{\phi}{M}\right)^{n^{N}_{ij}}M\bar{N}^{c}_{i}N_{j}+\text{h.c.},$ (2.1)
where $\{Q_{i},u_{i},d_{i},L_{i},l_{i},N_{i},H\}$ denote the left-handed quarks, the right-handed up-type quarks, the right-handed down-type quarks, the left-handed leptons, the right-handed charged leptons, the right-handed neutrinos, and the Standard Model (SM) Higgs doublet with $H^{c}=i\sigma_{2}H^{\ast}$, respectively. Here, a complex scalar field $\phi$, which is called a flavon field, is introduced. We assume that a mass of the right-handed neutrinos $M$ corresponds to a cut-off scale of the model to explain the light neutrino masses through the seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia . The Yukawa couplings $\{y^{u}_{ij},y^{d}_{ij},y^{l}_{ij},y^{\nu}_{ij},y^{N}_{ij}\}$ are ${\cal O}(1)$ real constants.
@@
[…]
@@
$f_{a}\equiv\frac{\sqrt{2}|v_{\phi}|}{N_{\mathrm{DW}}}=\frac{\sqrt{2}|\eta|M}{N_{\mathrm{DW}}}\,$, (2.20)
@@
For the phenomenological study, it is also useful to consider an effective coupling between the flaxion and the photon. The effective Lagrangian of the interaction is given by
@@ 2.2 Constraints from flavor physics
[…]
$f_{a}\gtrsim 2.5\times 10^{10}\,\mbox{GeV}\frac{30}{N_{\mathrm{DW}}}\left|\frac{(\kappa^{d}_{\rm AH})_{12}}{m_{s}-m_{d}}\right|.$ (2.27)
This bound can be evaluated by determining appropriate parameters of the FN model. The specific ranges of $\eta,M$ and $q$ to be examined in this study are described later, but in that domain, this inequality is almost automatically satisfied.
@@ 2.3 Constraints from cosmology
As discussed later, the misalignment angle is fixed through Eq. (2.29) for each model found by RL, so the upper bound of $H_{\rm inf}$ is given for each of the models.
@@ 3 Reinforcement learning
In RL, the subject of learning is called an agent, and the problem to be solved is called an environment. The learning process is constructed from three stages: observing the environment, choosing an action, and getting rewards. By repeating these procedures, it is known that the agent autonomously acquires a principle of actions that maximizes the sum of rewards. Details of RL and Q-learning, which are basic ingredients of RL, are explained in Ref. RL . In this paper, we adopt Deep Q-Network (DQN) as a RL algorithm, and use the package “gym” developed by OpenAI. Since the architecture in this work is similar to our previous work Nishimura:2023nre , we briefly review the RL, and setups of the learning will be explained by focusing on the key points. For more details about the architecture, see Ref. Nishimura:2023nre .
The flavor structure of quarks and leptons is determined by the charge shown in Eq. (2.2). In Ref. Nishimura:2023nre , many realistic models of the lepton sector are found in contrast to the quark sector. In other words, there are many varieties of the $U(1)$ charges of leptons leading to successful mass hierarchies and the flavor structure. On the other hand, viable solutions for the quark sector have proven difficult to obtain, even in the absence of renormalization effects. This fact raises concerns that the existence of renormalization effects will obstruct the agents from reaching valid solutions and make the learning of RL unstable. Moreover, the benchmarks established in the previous work, which neglected renormalization effects, are not directly applicable to the quark sector in a precise sense. Consequently, those solutions require modification to account for the influence of radiative corrections. Hence, at the moment, we will perform RL on the quark sector to clarify the effectiveness of RL and to compare the different energy scales.
@@ 3.1 The environment
with $i=1,2,3$, so the vector has 10 elements. This is the input data for neural networks which decide an action of the agent. To avoid a hierarchy of $U(1)$ charges, we take the value of each charge within:
@@
corresponding to total $19^{10}\,\sim\,10^{12}$ possibilities for only the quark sector. In a general $U(1)$ charge pair, non-integers $n_{ij}$ can be realized. To avoid this situation, we initialize the set of charges as $q(\phi)=+1$ with 50% probability and $q(\phi)=-1$ with 50% probability. This makes it easier to generate pairs of charges such that the denominator of Eq. (2.4) and Eq. (2.5) is $\pm 1$, so integers $n_{ij}$ are preferentially realized.
3.2 Neural Network
We utilize neural networks to determine an action of the agent. Their architecture is summarized in Table 1. The activation function is chosen as the SELU function for hidden layers and the softmax function for an output layer. We employ the ADAM optimizer in TensorFlow DBLP:journals/corr/AbadiABBCCCDDDG16 , and the Huber function is used as a loss function:
@@ 3.2 Neural Network
with $\delta=1$. This function combining the mean squared error and the mean absolute error will lead to the efficient learning by taking advantage of the features of both.
@@ 3.3 Agent
where $A$ corresponds to $\{Q_{i},u_{i},d_{i},\phi\}$. At an initial stage of the learning, for each environment, the ${\cal O}(1)$ coefficients in Yukawa terms Eq. (2.1) are picked up from the two Gaussian distribution with an average $\pm 1$ and a standard deviation 0.25. After the training with the neural networks introduced in the previous subsection, the coefficients are optimized to proper values by the Monte-Carlo simulation. Note that the coefficient $y^{u}_{33}$ was fixed at 0.4 in both the initialization and the optimization to ensure that the mass of the top quark is reproduced.
The agent chooses an action based on $\varepsilon$-greedy method. In this method, $\mathfrak{a}$ is determined by the neural network with probability $1-\varepsilon$ and randomly with probability $\varepsilon$ as follows:
$\mathfrak{a}=\left\{\begin{array}[]{l}\mathfrak{b}\quad(\text{with}\,1-\varepsilon)\\ \mathfrak{c}\quad(\text{with}\,\varepsilon)\end{array}\right$., (3.8)
where $\mathfrak{b}$ is called a greedy action, and $\mathfrak{c}$ represents a random action. By repeating choosing an action, a sequence of charges is defined as follows:
${\cal Q}^{\,\left(1\right)}\xrightarrow{\mathfrak{a}_{1}}{\cal Q}^{\,\left(2\right)}\xrightarrow{\mathfrak{a}_{2}}{\cal Q}^{\,\left(3\right)}\xrightarrow{\mathfrak{a}_{3}}\cdots,$ (3.9)
and this chain is called an episode. The initial set of charges ${\cal Q}^{\,\left(1\right)}$ is determined randomly. The number of actions is specified by $N_{\rm step}$ for one episode and the agent repeats the steps $N_{\rm ep}$ times. We took $\varepsilon$ as the following form to ensure that the agent gradually takes the greedy action:
$\varepsilon=\max\left(\varepsilon_{0}r^{k-1},\ \varepsilon_{\rm min}\right),$ (3.10)
with $k=1,2,...,N_{\rm ep}$. Following the same way as in Ref. Nishimura:2023nre , we adopt $\varepsilon_{0}=1,\ r=0.99999$ and $\varepsilon_{\rm min}=0.01$. Other hyper-parameters are set as $N_{\rm ep}=10^{5}$ for an episode number and $N_{\rm step}=32$ for a step number. Moreover, a batch size is 32, an epoch number is 32, and a learning rate is $\alpha=2.5\times 10^{-4}$, respectively  [^2] For simplicity, we use the same hyper-parameters as in the previous study, but the choice of hyper-parameters is also important as a factor that determines the success or failure of the learning. There are various optimization methods for hyper-parameters. Then, “Hyperopt” Bergstra_2015 and “Optuna” 10.1145/3292500.3330701 are known as libraries that provide tuning methods based on Bayesian optimization. In general, a “black box function” refers to a function whose functional form cannot be revealed in advance and whose value is determined only after the arguments are given. Bayesian optimization and genetic algorithms are effective methods for solving optimization problems for black box functions..
@@
    For the mixing angles, we focused on reproducing the characteristics of weak mixing rather than the exact experimental values. Following the previous studies utilizing RL, the ratios of the best-fit values are used to train the agent for the quark sector simply, and the uncertainties associated with the CKM matrix are not taken into account. In addition, only one flavon field is introduced in the Lagrangian Eq. (2.1). Thus, $\mathcal{C}$ does not include the evaluation of the CP phase.
@@
A large intrinsic value indicates that an obtained charge assignment well reproduces the renormalized masses and mixings  [^3] In the present study, the intrinsic value is defined the same as in Ref. Nishimura:2023nre to be based on the findings of the previous study. If the $\chi^{2}$ value is adopted for intrinsic value, the search can be conducted with more statistical significance.. Such charge assignment is called a terminal state. Specifically, the terminal state is defined to satisfy the following requirements:
@@
In this paper, we adopt $V_{0}=10.0$, $V_{1}=1.75$, and $V_{2}=0.2$. Here, $V_{1}=1.75$ ($V_{2}=0.2)$ means that a ratio of the calculated masses (mixings) to the renormalized masses (observed mixings) satisfies $1.78\times 10^{-2}\leq r_{\rm mass}\leq 56.2$ ($0.63\leq r_{\rm mixings}\leq 1.58$). The difference between $V_{1}$ and $V_{2}$ occurs from relative errors associated with the observed values of the masses and mixing matrices. While the error for the mass is ${\cal O}(1)$-${\cal O}(10)$%, the elements of the CKM matrix contain ${\cal O}(10^{-3})$-${\cal O}(1)$% error. Therefore, it is reasonable to search for a mixing matrix with a narrower tolerance than that of the masses. Note that the detailed values of $\{V_{0},V_{1},V_{2}\}$ are not completely determined by any specific principle, as these values are hyperparameters in our architecture.
We change the threshold value from $V_{1}=1.0$ used in Ref. Nishimura:2023nre to $V_{1}=1.75$. That work used experimental masses at the electroweak scale, which are shown in Particle Data Group ParticleDataGroup:2022pth . On the other hand, this study considers renormalized masses that are smaller than those used in the previous research. For example, the up-quark mass shown in Ref. ParticleDataGroup:2024prd is $m_{u,\mathrm{exp}}=2.16\,\mbox{MeV}$, and an allowed range for a terminal state in the case of $V_{1}=1.0$ is $0.216\,\mbox{MeV}\leq m_{u}\leq 21.6\,\mbox{MeV}$. However, $m_{u,\mathrm{RG}}=0.49\,\mbox{MeV}$ for $M=10^{14}\,\mbox{GeV}$ as shown in Table 2, so an allowed range with the same $V_{1}$ is $0.049\,\mbox{MeV}\leq m_{u}\leq 4.9\,\mbox{MeV}$. This variation is quantified as a tolerance $\Delta m_{u}$ in the following way:
@@
Since changes of the renormalized masses from $10^{14}\,\mbox{GeV}$ to $10^{17}\,\mbox{GeV}$ are relatively small, the same $V_{1}$ is adopted for all energy scales. Then, after optimizing the ${\cal O}(1)$ coefficients, we manually extract terminal states where $n_{ij}$ is positive. Moreover, at this time, the reproducibility of the masses is ensured by extracting models that satisfy $V_{1}=1.75\,(0.1\leq r_{\rm mass}\leq 10.0)$.
Let us denote the charge assignment $\cal Q$} as observed by the agent, and $\cal Q$$^{\prime}$ followed by the action $\mathfrak{a}$. For a pair of the assignment and the action $({\cal Q},\mathfrak{a})$, we give the agent rewards $\mathcal{R}$. The rewarding procedure is exactly the same as in Ref. Nishimura:2023nre in which the design of the rewards is described by:
    1. Give the basic point $\mathcal{R}_{\rm base}$, depending on the value of intrinsic value: $\mathcal{R}_{\rm base}=\left\{\begin{array}[]{ll}{\cal V}({\cal Q}^{\prime})-{\cal V}({\cal Q})&\qquad{\rm if}\,{\cal V}({\cal Q}^{\prime})-{\cal V}({\cal Q})>0\\ \mathcal{R}_{\rm offset}&\qquad{\rm if}\,{\cal V}({\cal Q}^{\prime})-{\cal V}({\cal Q})\leq 0\end{array}\right$., (3.23) where $\mathcal{R}_{\rm offset}$ corresponds to a penalty, chosen as $\mathcal{R}_{\rm offset}=-10$. 2. When the $\cal Q$$^{\prime}$ lies outside $9\leq{\cal Q}^{\prime}\leq 9$ or the flavon charge satisfies $q(\phi)=0$, we give the penalty ${\cal R}_{\rm offset}$ and the environment comes back to the original charge assignment $\cal Q$}. 3. When the $\cal Q$$^{\prime}$ is turned out to be a terminal state, we give the bonus point ${\cal R}_{\rm term}$, chosen as $\mathcal{R}_{\rm term}=100$. 4. Summing up the above points, we define the reward ${\cal R}({\cal Q},\mathfrak{a})$.
The success or failure of RL depends greatly on the design of the reward [^4] In many cases, the appropriate design of rewards requires heuristic construction depending on the environment and action space. In contrast, it has been pointed out that the design may be systematized by combining Bayesian statistics with RL levine2018reinforcementlearningcontrolprobabilistic Alternatively, by learning based on models in Bayesian statistics, various algorithms have been developed that can perform learning even with small amounts of data. “PILCO” 10.5555/3104482.3104541 is one example. We will leave these comprehensive studies for future work.. If the intrinsic value is drastically changed, it may become difficult to find FN charges. In order to reduce such uncertainties in our research, we do not include cosmological restrictions in the intrinsic value.